# Primordial neutrinos fade to gray: constraints from cosmological observables

### v1

@@ III BBN calculations
We include the modified neutrino distributions (8) in the BBN code PRIMAT Pitrou et al. (2018), extending the method used in Refs. Froustey and Pitrou (2020). In that work, in addition to the change in energy density due to distortions, the weak interaction rates of reactions (12) were modified at the Born approximation level (i.e., in the infinite nucleon mass approximation). Given the large distortions considered here, we include the effect of spectral distortions consistently with the various corrections to the weak rates derived in Pitrou et al. (2018), namely, radiative and finite nucleon mass corrections.
@@
This picture is confirmed in Fig. 4, where we plot the PRIMAT results for the helium-4 ($Y_{\mathrm{p}}$) and deuterium ($\mathrm{D/H}$) abundances, setting the baryon abundance value to $\Omega_{b}h^{2}=0.0224$. We also show the experimental measurements with colored bands, see Sec. IV for details.
@@ IV Constraints from observations
Given their effects on cosmological observables (see Sec. II.2), the two parameters $\{y_{g},\Delta N\}$ can be constrained with observations. Here, we will use observations of primordial abundances from BBN and constraints on the cosmological $N_{\mathrm{eff}}$ parameter coming from non-BBN observations, namely the cosmic microwave background, CMB lensing and baryon acoustic oscillations (BAO). More specifically, we use CMB and CMB lensing observations from Planck Aghanim et al. (2020a, b), and BAO observations from DESI Abdul Karim et al. (2025).
@@
One could also consider the lithium-7 abundance, whose recommended value by the PDG Navas et al. (2024) (using observational data from Sbordone et al. (2010)) is $\mathrm{Li/H}^{\mathrm{obs}}=(1.6\pm 0.3)\times 10^{-10}$. The lithium observations are known to be in strong tension with predictions from standard BBN theory—the so-called “lithium problem” Fields (2011). Although one could hope that a combination of our non-standard parameters could satisfy the observations (14)–(15) and provide a lithium abundance compatible with observations, this is not the case. The parameter values required for the observed lithium abundance are in large tension with the parameter space preferred for helium and deuterium measurements. Given the uncertain status of the lithium problem Navas et al. (2024), we leave the lithium abundance out of the BBN observables used for our constraints.
@@
To calculate the parameter constraints, we employ the Markov chain Monte Carlo (MCMC) code Cobaya Torrado and Lewis (2021) and its MCMC sampler Lewis and Bridle (2002); Lewis (2013); Neal (2005). The inference of cosmological parameters from observations in Cobaya requires likelihoods associated to BBN and CMB + lensing + BAO measurements. For BBN observations, we employ Gaussian likelihoods around the values of the measured abundances. To calculate the predicted abundances for a given set of parameters ($y_{g}$, $\Delta N$ and baryon abundance $\omega_{b}\equiv\Omega_{b}h^{2}$), we construct a table with the code PRIMAT (see Sec. III). To get the standard deviation of the Gaussian likelihoods, we quadratically add the observational error and the error of PRIMAT’s predictions due to nuclear rates uncertainties. We use the code CLASS Blas et al. (2011) to calculate the other cosmological observables (CMB, lensing, BAO). There are 8 sampled parameters: the 6 standard $\Lambda$CDM cosmological parameters, along with $\Delta N$ and $y_{g}$; $Y_{\mathrm{p}}$ is not left to vary but is obtained from the PRIMAT table. We thus have to modify CLASS at two levels: first, to include the effect of the distorted distributions (8) (instead of the pure Fermi-Dirac ones), then, to read our BBN table, which has one more parameter than CLASS standard BBN tables because of the presence of the $y_{g}$ parameter. We then use the CamSpec Rosenberg et al. (2022) likelihood, which makes use of Planck’s PR4 (NPIPE) Akrami et al. (2020) maps.
@@
In principle, we can also constrain the baryon abundance, $\omega_{b}=\Omega_{b}h^{2}$, using BBN. For BBN constraints, we use a prior on this abundance obtained with a run including the aforementioned CMB, CMB lensing and BAO observations, but without using any BBN information. In practice, we find that these observations are much more constraining on $\omega_{b}$ than BBN, so the $\omega_{b}$ constraints from the prior remain essentially unaltered.
@@ IV.1 Baseline results
Figure 5 shows the constraints in the $(y_{g},\Delta N)$ plane (marginalized over $\omega_{b}$) coming from the aforementioned observations: deuterium, helium, and CMB + lensing + BAO. The figure shows all 3 constraints separately, as well as the joint constraint that takes them all into account. Note that all three constraints contain a certain degeneracy between $\Delta N$ and $y_{g}$ that they cannot individually break. The main degeneracy is due to the equivalent gravitational effects of the $y_{g}$-distortion and some extra radiation: comparison of Eqs. (9) and (11) shows that the extra energy density of three flavors of neutrinos and antineutrinos distorted with the same parameter $y_{g}$ can be compensated by $\Delta N\simeq-3y_{g}$. This leads to the “diagonal” degeneracy on Fig. 5, which is notably broken by the strongly nonlinear dependence of the helium abundance on $y_{g}$ (see Fig. 4). This dependence gives rise to the curved shape of the helium constraint in Fig. 5. The slight sensitivity to $Y_{\mathrm{p}}$ of CMB measurements also explains why the red contours on Fig. 5 are not perfectly along the $\Delta N\simeq-3y_{g}$ line.
@@ IV.2 Alternative nuclear rates
Comparing Figs. 4 and 6, one can see, as expected, that the helium abundance is almost unchanged, while the deuterium predictions shift to higher values, reconciling the measurement (15) and the prediction for $\mathrm{D/H}$ at $(y_{g},\Delta N)=(0,0)$.
@@ IV.3 Alternative helium observations
We report this value as a purple band on the left panels of Figs. 4 and 6 ; it is $1.7\sigma$ below the traditional value, consistent with many observations and recommended by the PDG Navas et al. (2024). Such a smaller value can be accommodated by a nonzero lepton asymmetry in the electron neutrino sector at the epoch of BBN Matsumoto et al. (2022); Burns et al. (2023); Escudero et al. (2023); Froustey and Pitrou (2024), i.e., a neutrino $\mu$-distortion (see Sec. II). Here, we check in the following how our results change using the measurement (18) instead of (14), exploring the potential of these generalized $y$-distortions to accommodate this helium abundance.

### v2

@@ III BBN calculations
We include the modified neutrino distributions (8) in the BBN code PRIMAT Pitrou et al. (2018), extending the method used in Ref. Froustey and Pitrou (2020). In that work, in addition to the change in energy density due to distortions, the weak interaction rates of reactions (12) were modified at the Born approximation level (i.e., in the infinite nucleon mass approximation). Given the large distortions considered here, we include the effect of spectral distortions consistently with the various corrections to the weak rates derived in Pitrou et al. (2018), namely, radiative and finite nucleon mass corrections.
@@
This picture is confirmed in Fig. 6, where we plot the PRIMAT results for the helium-4 ($Y_{\mathrm{p}}$) and deuterium ($\mathrm{D/H}$) abundances, setting the baryon abundance value to $\Omega_{b}h^{2}=0.0224$ (preferred value from Planck Aghanim et al. (2020b)). We also show the experimental measurements with colored bands, see Sec. IV for details.
@@ IV Constraints from observations
Given their effects on cosmological observables (see Sec. II.2), the two parameters $\{y_{g},\Delta N\}$ can be constrained with observations. Here, we will use observations of primordial abundances from BBN and constraints on the cosmological $N_{\mathrm{eff}}$ parameter coming from non-BBN observations, namely the cosmic microwave background, CMB lensing and baryon acoustic oscillations (BAO). More specifically, we use CMB and CMB lensing observations from Planck Aghanim et al. (2020a, c), and BAO observations from DESI Abdul Karim et al. (2025).
@@
One could also consider the lithium-7 abundance, whose recommended value by the Particle Data Group Navas et al. (2024) (using observational data from Sbordone et al. (2010)) is $\mathrm{Li/H}^{\mathrm{obs}}=(1.6\pm 0.3)\times 10^{-10}$. The lithium observations are known to be in strong tension with predictions from standard BBN theory—the so-called “lithium problem” Fields (2011). Although one could hope that a combination of our non-standard parameters could satisfy the observations (14)–(15) and provide a lithium abundance compatible with observations, this is not the case. The parameter values required for the observed lithium abundance are in large tension with the parameter space preferred for helium and deuterium measurements. Given the uncertain status of the lithium problem Navas et al. (2024), we leave the lithium abundance out of the BBN observables used for our constraints.
@@
To calculate the parameter constraints, we employ the Markov chain Monte Carlo (MCMC) code Cobaya Torrado and Lewis (2021) and its MCMC sampler Lewis and Bridle (2002); Lewis (2013); Neal (2005). The inference of cosmological parameters from observations in Cobaya requires likelihoods associated to BBN and CMB + lensing + BAO measurements. For BBN observations, we employ Gaussian likelihoods around the values of the measured abundances. To calculate the predicted abundances for a given set of parameters ($y_{g}$, $\Delta N$ and baryon abundance $\omega_{b}\equiv\Omega_{b}h^{2}$), we construct a table with the code PRIMAT (see Sec. III). To get the standard deviation of the Gaussian likelihoods, we quadratically add the observational error and the error of PRIMAT’s predictions due to nuclear rates uncertainties. We use the code CLASS Blas et al. (2011) to calculate the other cosmological observables (CMB, lensing, BAO). There are eight sampled parameters: the six standard $\Lambda$CDM cosmological parameters, along with $\Delta N$ and $y_{g}$; $Y_{\mathrm{p}}$ is not left to vary but is obtained from the PRIMAT table. We thus have to modify CLASS at two levels: first, to include the effect of the distorted distributions (8) (instead of the pure Fermi-Dirac ones)  [^7] In the results that we show, to be able to consider the case $\Delta N<0$, we implement the distorsions in CLASS in an effective way that only considers the effect on the energy density and the helium fraction, setting $N_{\mathrm{eff}}$ and $Y_{\mathrm{p}}$ to the corresponding value given $y_{g}$ and $\Delta N$. In the regime $\Delta N\geq 0$, we have checked that this is consistent with what we get if we implement the full physical distribution function in CLASS., then, to read our BBN table, which has one more parameter than CLASS standard BBN tables because of the presence of the $y_{g}$ parameter. We then use the CamSpec Rosenberg et al. (2022) likelihood, which makes use of Planck’s PR4 (NPIPE) Akrami et al. (2020) maps.
@@
In principle, we can also constrain the baryon abundance, $\omega_{b}=\Omega_{b}h^{2}$, using BBN. For BBN-only constraints, we use a prior on this abundance obtained with a run including the aforementioned CMB, CMB lensing and BAO observations, but without using any BBN information. More specifically, this is a Gaussian prior with $\omega_{b}=0.0223979\pm 0.0001675$. In practice, we find that these observations are much more constraining on $\omega_{b}$ than BBN, so the $\omega_{b}$ constraints from the prior remain essentially unaltered.
@@ IV.1 Baseline results
Figure 7 shows the constraints in the $(y_{g},\Delta N)$ plane (marginalized over $\omega_{b}$) coming from the aforementioned observations: deuterium, helium, and CMB + lensing + BAO. The figure shows all 3 constraints separately, as well as the joint constraint that takes them all into account. Note that all three constraints contain a certain degeneracy between $\Delta N$ and $y_{g}$ that they cannot individually break. The main degeneracy is due to the equivalent gravitational effects of the $y_{g}$-distortion and some extra radiation, as discussed in Fig. 4. This leads to the “diagonal” degeneracy on Fig. 7, which is notably broken by the strongly nonlinear dependence of the helium abundance on $y_{g}$ (see Fig. 6). This dependence gives rise to the curved shape of the helium constraint in Fig. 7. The slight sensitivity to $Y_{\mathrm{p}}$ of CMB measurements also explains why the red contours on Fig. 7 are not perfectly along the $\Delta N\simeq-3y_{g}$ line.
@@ IV.2 Alternative nuclear rates
Comparing Figs. 6 and 8, one can see, as expected, that the helium abundance is almost unchanged, while the deuterium predictions shift to higher values, reconciling the measurement (15) and the prediction for $\mathrm{D/H}$ at $(y_{g},\Delta N)=(0,0)$.
@@ IV.3 Alternative helium observations
We report this value as a purple band on the left panels of Figs. 6 and 8 ; it is $1.7\sigma$ below the traditional value, consistent with many observations and recommended by the Particle Data Group Navas et al. (2024). Such a smaller value can be accommodated by a nonzero lepton asymmetry in the electron neutrino sector at the epoch of BBN Matsumoto et al. (2022); Burns et al. (2023); Escudero et al. (2023); Froustey and Pitrou (2024), i.e., a neutrino $\mu$-distortion (see Sec. II). Here, we check in the following how our results change using the measurement (18) instead of (14), exploring the potential of these generalized $y$-distortions to accommodate this helium abundance.